# Class-Based Expurgation Attains Csiszár’s Expurgated Source-Channel Exponent

### v1

@@ III-C Two-Class Partition of the Source Sequences
An immediate implication of Corollary 3 is that $E_{J,2}^{\mathrm{ex}}(t,P_{V})$ can be achieved with only two classes, in contrast to Csiszár’s construction [4], which requires $m=N_{k}$ classes.
IV Proofs
Due to space constraints, we present only the direct derivations of the dual-domain results, as they are less straightforward and not as immediate as those in the primal domain. The primal-domain case follows by similar arguments, using standard tools from the method of types.
IV-A Proof of Lemma 3
The proof closely follows the approach in [5, Sec. III.C]. We consider the code from Lemma 1, and use the following decoding metric
$q(\boldsymbol{v},\boldsymbol{x},\boldsymbol{y})=W^{n}(\boldsymbol{y}|\boldsymbol{x})\exp\!\big{(}2kD(P_{i}\|P_{V})\big{)}$ (25)
if $\boldsymbol{v}\in\mathcal{T}^{k}(P_{i})$. Then, for any $s>0$, we can upper bound $p_{e}\big{(}\boldsymbol{v},j,\mathsf{C}\big{)}$ as
$\!\!p_{e}\big{(}\boldsymbol{v},j,\mathsf{C}\big{)}\!$ $=\mathbb{P}\!\left[\bigcup_{\bar{\boldsymbol{v}}\in\mathcal{T}^{k}(P_{j})}\!\!\!q(\boldsymbol{\bar{v}},\boldsymbol{x}_{\bar{\boldsymbol{v}}},\boldsymbol{Y})\geq q(\boldsymbol{v},\boldsymbol{x}_{\boldsymbol{v}},\boldsymbol{Y})\right]$ (26)
$=\mathbb{P}\!\left[\bigcup_{\bar{\boldsymbol{v}}\in\mathcal{T}^{k}(P_{j})}\!\!\!q(\boldsymbol{\bar{v}},\boldsymbol{x}_{\bar{\boldsymbol{v}}},\boldsymbol{Y})^{s}\geq q(\boldsymbol{v},\boldsymbol{x}_{\boldsymbol{v}},\boldsymbol{Y})^{s}\right]$ (27)
$\leq\sum_{\boldsymbol{y}}W^{n}(\boldsymbol{y}|\boldsymbol{x}_{\boldsymbol{v}})\!\!\sum_{\bar{\boldsymbol{v}}\in\mathcal{T}^{k}(P_{j})}\left(\frac{q(\boldsymbol{\bar{v}},\boldsymbol{x}_{\bar{\boldsymbol{v}}},\boldsymbol{y})}{q(\boldsymbol{v},\boldsymbol{x}_{\boldsymbol{v}},\boldsymbol{y})}\right)^{s}.$ (28)
Substituting the decoding metric and setting $s=\frac{1}{2}$, we obtain
$\begin{split}p_{e}\big{(}\boldsymbol{v},j,\mathsf{C}\big{)}\,&\leq\left(\frac{\exp\left(kD\left(P_{j}\|P_{V}\right)\right)}{\exp\left(kD\left(P_{i}\|P_{V}\right)\right)}\right)\\ &\times\sum_{\bar{\boldsymbol{v}}\in\mathcal{T}^{k}(P_{j})}\sum_{\boldsymbol{y}}\sqrt{W^{n}(\boldsymbol{y}|\boldsymbol{x}_{\boldsymbol{v}})W^{n}(\boldsymbol{y}|\boldsymbol{x}_{\bar{\boldsymbol{v}}})}.\end{split}$ (29)
If we assume $\rho_{ij}\geq 1$ for all $(i,j)$, we can use Hölder’s inequality [6, Problem 4.15.f] to move the summation over $\boldsymbol{\bar{v}}$ outside, as follows
$\begin{split}\!\!p_{e}\big{(}\boldsymbol{v},j,\mathsf{C}\big{)}^{\frac{1}{\rho_{ij}}}\!&\leq\left(\frac{\exp(kD(P_{j}\|P_{V}))}{\exp(kD(P_{i}\|P_{V}))}\right)^{\frac{1}{\rho_{ij}}}\\ &\times\!\!\!\!\!\sum_{\bar{\boldsymbol{v}}\in\mathcal{T}(P_{j})}\prod_{l=1}^{n}\!\left[\sum_{y}\!\sqrt{W(y|x_{\boldsymbol{v}}^{l})W(y|x_{\boldsymbol{\bar{v}}}^{l})}\right]^{\frac{1}{\rho_{ij}}}\!\!\!,\end{split}$ (30)
where $x_{\boldsymbol{v}}^{l}$ denotes the $l$-th element of $\boldsymbol{x}_{\boldsymbol{v}}$. We now take the expectation of the above expression
$\begin{split}\mathbb{E}\left[p_{e}\big{(}\boldsymbol{v},j,\mathsf{C}\big{)}^{\frac{1}{\rho_{ij}}}\right]\,&\leq\left(\frac{\exp(kD(P_{j}\|P_{V}))}{\exp(kD(P_{i}\|P_{V}))}\right)^{\frac{1}{\rho_{ij}}}\\ &\times\mathbb{E}\left[\sum_{\bar{\boldsymbol{v}}\in\mathcal{T}^{k}(P_{j})}\prod_{l=1}^{n}\,e^{\frac{-d_{B}(X_{l},\bar{X}_{l})}{\rho_{ij}}}\right],\end{split}$ (31)
where the expectation is taken with respect to $P_{\boldsymbol{X}}^{(\hat{i})}\times P_{\boldsymbol{X}}^{(\hat{j})}$, where $P_{\boldsymbol{X}}^{(c)}$ denotes the uniform distribution over the type class $\mathcal{T}^{n}(Q_{c})$. Here, $\hat{i}$ denotes the index of the class to which source type $i$ belongs. We can evaluate this expectation as follows
$\begin{split}&\mathbb{E}\left[\,\prod_{l=1}^{n}\,e^{\frac{-d_{B}(X_{l},\bar{X}_{l})}{\rho_{ij}}}\right]\\ &\hskip 28.45274pt=\sum_{(\boldsymbol{x},\bar{\boldsymbol{x}})}P_{\boldsymbol{X}}^{(\hat{j})}({\boldsymbol{x}})P_{\boldsymbol{X}}^{(\hat{i})}(\bar{\boldsymbol{x}})\prod_{k=1}^{n}\,e^{\frac{-d_{B}(x_{l},\bar{x}_{l})}{\rho_{ij}}}.\end{split}$ (32)
We use Lemma 2.1 of [12] to rewrite the first expectation as if the distribution were i.i.d. with $Q_{\hat{i}}$. This gives
$\begin{split}&\mathbb{E}\left[\,\prod_{l=1}^{n}\,e^{\frac{-d_{B}(X_{l},\bar{X}_{l})}{\rho_{ij}}}\right]\\ &\hskip 14.22636pt\,\dot{\leq}\sum_{\boldsymbol{x}\in\mathcal{T}^{n}(Q_{\hat{j}})}\frac{1}{\big{|}\mathcal{T}^{n}(Q_{\hat{j}})\big{|}}\prod_{l=1}^{n}\sum_{\bar{x}}Q_{\hat{i}}(x)\,e^{\frac{-d_{B}(x_{l},\bar{x})}{\rho_{ij}}}\end{split}$ (33)
$\hskip 14.22636pt\doteq\,\prod_{x\in\mathcal{X}}\left[\sum_{\bar{x}}Q_{\hat{i}}(x)\,e^{\frac{-d_{B}(x,\bar{x})}{\rho_{ij}}}\right]^{nQ_{\hat{j}}(x)}$ (34)
$\hskip 14.22636pt=\exp\left(n\sum_{x}Q_{\hat{j}}(x)\log\sum_{\bar{x}}Q_{\hat{i}}(x)\,e^{\frac{-d_{B}(x,\bar{x})}{\rho_{ij}}}\right).$ (35)
Note that the bound in (35) does not depend on $\boldsymbol{\bar{v}}$, and so
$\begin{split}&\,\mathbb{E}\left[\sum_{\bar{\boldsymbol{v}}\in\mathcal{T}^{k}(P_{j})}\prod_{k=1}^{n}\,e^{\frac{-d_{B}(X_{k},\bar{X}_{k})}{\rho_{ij}}}\right]\\ &\,\,\dot{\leq}\exp\!\left(-n\!\left[-\sum_{x}Q_{\hat{j}}(x)\log\sum_{\bar{x}}Q_{\hat{i}}(x)\,e^{\frac{-d_{B}(x,\bar{x})}{\rho_{ij}}}\!\!-R_{j}\right]\right).\end{split}$ (36)
Raising both sides to the power $\rho_{ij}$ and taking the worst-case choice of $Q_{\hat{i}}$, we obtain
$\mathbb{E}\left[p_{e}\big{(}\boldsymbol{v},j,\mathsf{C}\big{)}^{\frac{1}{\rho_{ij}}}\right]^{\rho_{ij}}$
$\leq\!\left(\frac{\exp(kD(P_{j}\|P_{V}))}{\exp(kD(P_{i}\|P_{V}))}\right)\exp\!\left(\!-n\!\left[E_{\mathrm{x}}^{\prime}({Q_{\hat{j}}},\rho_{ij})\!-\!\rho_{ij}R_{j}\right]\right).$ (37)
Having evaluated the expectation in (11), we now obtain the following bound on the error probability of the code in Lemma 1 as follows
$p_{e}(\mathcal{C})$ $=\sum_{\boldsymbol{v}}P^{k}(\boldsymbol{v})\,p_{e}(\boldsymbol{v},\mathcal{C})$ (38)
˙$\begin{split}&\,\,\dot{\leq}\,\,\sum_{(i,j)}\sum_{\boldsymbol{v}\in\mathcal{T}^{k}(P_{i})}P^{k}(\boldsymbol{v})\left(\frac{\exp(kD(P_{j}\|P_{V}))}{\exp(kD(P_{i}\|P_{V}))}\right)\\ &\hskip 28.45274pt\times\exp\left(-n\left[E_{\mathrm{x}}^{\prime}({Q_{\hat{j}}},\rho_{ij})-\rho_{ij}R_{j}\right]\right).\end{split}$ (39)
Observe that $\sum_{\boldsymbol{v}\in\mathcal{T}^{k}(P_{i})}P^{k}(\boldsymbol{v})\doteq\exp(kD(P_{i}\|P_{V}))$. This implies that
$\!\!\!\!\!\!\sum_{\boldsymbol{v}\in\mathcal{T}^{k}(P_{i})}\!\!\!\!\!P^{k}(\boldsymbol{v})\!\left(\frac{\exp(kD(P_{j}\|P_{V}))}{\exp(kD(P_{i}\|P_{V}))}\right)$ $\!\doteq\exp(kD(P_{j}\|P_{V})).$ (40)
We can further simplify this for any $\lambda_{j}>0$ as follows
$\begin{split}&\exp\big{(}kD(P_{j}\|P_{V})\big{)}\\ &\hskip 7.11317pt=\exp\left(k\sum_{v}P_{j}(v)\log\frac{P_{V}(v)}{P_{j}(v)}\right)\end{split}$ (41)
$\hskip 7.11317pt=\exp\!\left(k(1+\lambda_{j})\sum_{v}P_{j}(v)\log\frac{P_{V}(v)^{\frac{1}{1+\lambda_{j}}}}{P_{j}(v)P_{j}(v)^{\frac{-\lambda_{j}}{1+\lambda_{j}}}}\right).$ (42)
The last expression can be written as
$\begin{split}&\exp\left(k(1\!+\!\lambda_{j})\sum_{v}P_{j}(v)\log\frac{P_{V}(v)^{\frac{1}{1+\lambda_{j}}}}{P_{j}(v)P_{j}(v)^{\frac{-\lambda_{j}}{1+\lambda_{j}}}}\right)\\ &=\exp\!\left(\!n\!\left[t(1\!+\!\lambda_{j})\!\sum_{v}P_{j}(v)\log\frac{P_{V}(v)^{\frac{1}{1+\lambda_{j}}}}{P_{j}(v)}\!-\!\lambda_{j}R_{j}\right]\right)\end{split}$ (43)
$\leq\exp\!\left(\!n\!\left[t(1\!+\!\lambda_{j})\!\log\sum_{v}P_{j}(v)\frac{P_{V}(v)^{\frac{1}{1+\lambda_{j}}}}{P_{j}(v)}\!-\!\lambda_{j}R_{j}\right]\right)$ (44)
$=\exp\left(-n\Big{[}\lambda_{j}R_{j}\!-\!tE_{s}(\lambda_{j},P_{V})\Big{]}\right),$ (45)
where we applied Jensen’s inequality in (44). By setting $\rho_{ij}=\rho_{j}$, we can upper bound the expression in (39) as
$\begin{split}&p_{e}(\mathcal{C})\,\,\dot{\leq}\,\,\sum_{j=1}^{N_{k}}\exp\bigg{(}-n\Big{[}{E}_{\mathrm{x}}^{\prime}\big{(}{Q}_{\hat{j}},\rho_{j}\big{)}\\ &\hskip 56.9055pt+\big{(}\lambda_{j}-\rho_{j}\big{)}{R}_{j}-t\mathrm{E}_{\mathrm{s}}\big{(}\lambda_{j},P_{V}\big{)}\Big{]}\bigg{)}\end{split}$ (46)
$\begin{split}&\hskip 28.45274pt=\sum_{c=1}^{m}\sum_{i\in\Lambda_{c}}\exp\bigg{(}-n\Big{[}{E}_{\mathrm{x}}^{\prime}\big{(}{Q}_{c},\rho_{ic}\big{)}\\ &\hskip 56.9055pt+\big{(}\lambda_{ic}-\rho_{ic}\big{)}{R}_{i}-t\mathrm{E}_{\mathrm{s}}\big{(}\lambda_{ic},P_{V}\big{)}\Big{]}\bigg{)}.\end{split}$ (47)
To prove Corollary 1, observe that by applying (40), the bound in (39) can be rewritten as
$p_{e}(\mathcal{C})$$\,\dot{\leq}\,\sum_{i=1}^{N_{k}}\sum_{\boldsymbol{v}\in\mathcal{T}^{k}(P_{i})}\frac{P^{k}(\boldsymbol{v})}{\hat{P}_{\boldsymbol{v}}(\boldsymbol{v})^{\rho_{i}}}\exp\big{(}-n\,{E}_{\mathrm{x}}^{\prime}(Q_{\hat{i}},\rho_{i})\big{)}$ (48)
$=\sum_{c=1}^{m}\left(\sum_{\boldsymbol{v}\in\mathcal{A}_{c}}\frac{P^{k}(\boldsymbol{v})}{\hat{P}_{\boldsymbol{v}}(\boldsymbol{v})^{\rho_{c}}}\right)\exp\!\big{(}\!-\!n\,{E}_{\mathrm{x}}^{\prime}(Q_{c},\rho_{c})\big{)}.$ (49)
The inner sum can be upper bounded as
$\!\!\!\!\sum_{\boldsymbol{v}\in\mathcal{A}_{c}}\frac{P^{k}(\boldsymbol{v})}{\hat{P}_{\boldsymbol{v}}(\boldsymbol{v})^{\rho}}$ $=\sum_{i\in\Lambda_{c}}\sum_{\boldsymbol{v}\in\mathcal{T}^{k}(P_{i})}\frac{P^{k}(\boldsymbol{v})}{P_{i}^{k}(\boldsymbol{v})^{\rho}}$ (50)
$\doteq\sum_{i\in\Lambda_{c}}\exp\left(k\sum_{v}P_{i}(v)\log\frac{P_{V}(v)}{P_{i}(v)^{1+\rho}}\right)$ (51)
$=\sum_{i\in\Lambda_{c}}\!\left[\exp\!\left(\!k\!\sum_{v}\!P_{i}(v)\!\log\!\frac{P_{V}(v)^{\frac{1}{1+\rho}}}{P_{i}(v)}\!\right)\!\right]^{1+\rho}$ (52)
$=\sum_{i\in\Lambda_{c}}\left[\sum_{\boldsymbol{v}\in\mathcal{A}_{c}}P^{k}(\boldsymbol{v})^{\frac{1}{1+\rho}}\right]^{1+\rho}$ (53)
$\leq\left[\sum_{i\in\Lambda_{c}}\sum_{\boldsymbol{v}\in\mathcal{A}_{c}}P^{k}(\boldsymbol{v})^{\frac{1}{1+\rho}}\right]^{1+\rho}$ (54)
$=\left[\sum_{\boldsymbol{v}\in\mathcal{A}_{c}}P^{k}(\boldsymbol{v})^{\frac{1}{1+\rho}}\right]^{1+\rho},$ (55)
where in (51), we used the fact that $|\mathcal{T}^{k}(P_{i})|\doteq\exp(kH(P_{i}))$, and in (54), we applied the inequality $\sum_{i}a_{i}^{r}\leq(\sum_{i}a_{i})^{r}$ for $r\geq 1$ [6, Problem 4.15.f].
IV-B Proof of Theorem 2
We first set $\rho_{ic}=\rho_{c}$ and $\lambda_{ic}=\lambda_{c}$. To find the optimal partitioning, assume each class is associated with given values of $\rho_{c}$ and $\lambda_{c}$, which will be optimized later. The source type class $\mathcal{T}^{k}(P_{i})$ is then assigned to class $c^{\star}$ if
$\!c^{\star}\!=\!\underset{c\in\{1,\ldots,m\}}{\operatorname{argmax}}E_{\mathrm{x}}^{\prime}(\mathrm{Q}_{c},\rho_{c})\!+\!(\lambda_{c}\!-\!\rho_{c}){R}_{i}-t{E}_{\mathrm{s}}(\lambda_{c},P_{V}).$ (56)
This means that
$\begin{split}&p_{e}(\mathcal{C})\,\,\dot{\leq}\,\,\sum_{j=1}^{N_{k}}\exp\bigg{(}-n\bigg{[}\max_{c}\,{E}_{\mathrm{x}}^{\prime}\big{(}{Q}_{c},\rho_{c}\big{)}\\ &\hskip 56.9055pt+\big{(}\lambda_{c}-\rho_{c}\big{)}{R}_{j}-t\mathrm{E}_{\mathrm{s}}\big{(}\lambda_{c},P_{V}\big{)}\bigg{]}\bigg{)}\end{split}$ (57)
˙$\begin{split}&\hskip 24.18501pt\,\,\dot{\leq}\,\exp\bigg{(}-n\bigg{[}\min_{R}\max_{c}\,{E}_{\mathrm{x}}^{\prime}\big{(}{Q}_{c},\rho_{c}\big{)}\\ &\hskip 56.9055pt+\big{(}\lambda_{c}-\rho_{c}\big{)}{R}-t\mathrm{E}_{\mathrm{s}}\big{(}\lambda_{c},P_{V}\big{)}\bigg{]}\bigg{)}.\end{split}$ (58)
Optimizing over the parameters yields
$\begin{split}&p_{e}(\mathcal{C})\,\dot{\leq}\,\exp\bigg{(}-n\bigg{[}\min_{R}\max_{\lambda>0}\max_{\rho\geq 1}\max_{Q\in\mathcal{Q}_{m}}\,{E}_{\mathrm{x}}^{\prime}\big{(}{Q},\rho\big{)}\\ &\hskip 56.9055pt+\big{(}\lambda-\rho\big{)}{R}-t{E}_{\mathrm{s}}\big{(}\lambda,P_{V}\big{)}\bigg{]}\bigg{)}.\end{split}$ (59)
Since $E_{s}(\lambda,P_{V})$ is convex in $\lambda$, we can apply Sion’s minimax theorem [13] to exchange the order of $\min_{R}$ and $\max_{\lambda}$, which gives
$\begin{split}&p_{e}(\mathcal{C})\,\dot{\leq}\,\exp\bigg{(}-n\bigg{[}\max_{\lambda>0}\min_{R}\max_{\rho\geq 1}\max_{Q\in\mathcal{Q}_{m}}\,{E}_{\mathrm{x}}^{\prime}\big{(}{Q},\rho\big{)}\\ &\hskip 71.13188pt+\big{(}\lambda-\rho\big{)}{R}-t{E}_{\mathrm{s}}\big{(}\lambda,P_{V}\big{)}\bigg{]}\bigg{)}.\end{split}$ (60)
Observe that if $\lambda<1$, then $(\lambda-\rho)R<0$, which implies that ${R}\rightarrow\infty$. Hence, the range of $\lambda$ is restricted to $\lambda\geq 1$.
IV-C Proof of Theorem 3
Let $g(\lambda)\triangleq E_{\mathrm{x}}^{\prime}(\lambda,\mathcal{Q}_{m})$. Adopting from [2], the concave conjugate of $g(\cdot)$ over $[1,\infty)$ is given by [11]
$g^{\star}(\lambda)=\inf_{\rho\geq 1}\Big{\{}\lambda\rho-E_{\mathrm{x}}^{\prime}(\rho,\mathcal{Q}_{m})\Big{\}}.$ (61)
The concave hull of $g(\cdot)$ is given by its biconjugate $g^{\star\star}(\lambda)$. Hence, we obtain
$g^{\star\star}(\lambda)$ $=\inf_{R}\bigg{\{}\lambda R-\inf_{\rho\geq 1}\Big{\{}\rho R-E_{\mathrm{x}}^{\prime}(\rho,\mathcal{Q}_{m})\Big{\}}\bigg{\}}$ (62)
$=\inf_{R}\sup_{\rho\geq 1}\Big{\{}E_{\mathrm{x}}^{\prime}(\rho,\mathcal{Q}_{m})+(\lambda-\rho)R\Big{\}}.$ (63)
This argument closely follows the proof of [2, Lemma 3].
IV-D Proof of Lemma 4
Assume the two classes are associated with given parameters $\lambda_{c}$ and $\rho_{c}$. If $\boldsymbol{v}\in\mathcal{T}^{k}(P_{i})$ is assigned to class 1, then by (56), we have
$\begin{split}&E_{\mathrm{x}}^{\prime}({Q}_{1},\rho_{1})+(\lambda_{1}-\rho_{1}){R}_{i}-t{E}_{\mathrm{s}}(\lambda_{1},P_{V})\\ &\hskip 21.33955pt\geq E_{\mathrm{x}}^{\prime}({Q}_{2},\rho_{2})+(\lambda_{2}-\rho_{2}){R}_{i}-t{E}_{\mathrm{s}}(\lambda_{2},P_{V}).\end{split}$ (64)
Rearranging terms, this implies
$\!R_{i}\!\leq\!\frac{E_{\mathrm{x}}^{\prime}({Q}_{1},\rho_{1})\!-\!E_{\mathrm{x}}^{\prime}({Q}_{2},\rho_{2})\!+\!t\left[{E}_{\mathrm{s}}(\lambda_{2},P_{V})\!-\!t{E}_{\mathrm{s}}(\lambda_{1},P_{V})\right]}{\big{(}\lambda_{2}-\rho_{2}\big{)}-\big{(}\lambda_{1}-\rho_{1}\big{)}}.$ (65)
Similarly, if $\boldsymbol{v}\in\mathcal{T}^{k}(P_{i})$ is assigned to class 2, the inequality in (65) is reversed.

### v2

@@ III-C Two-Class Partition of the Source Sequences
An immediate implication of Corollary 3 is that $E_{J,2}^{\mathrm{ex}}(t,P_{V})$ can be achieved with only two classes, in contrast to Csiszár’s construction [4], which requires $m=N_{k}$ classes, with $N_{k}$ growing polynomially in $k$.
IV Single Class Coding
This section presents the dual-domain counterparts of our results in the absence of partitioning, where all codewords share the same composition. The following lemmas present the counterparts of Theorem 2 and Corollary 2 in this setting
Lemma 5.
Assume that all codewords are drawn independently and uniformly from the type class $\mathcal{T}^{n}(Q)$, for some given $Q\in\mathcal{P}_{n}(\mathcal{X})$. Then, there exists a codebook such that
$p_{e}\,\,\dot{\leq}\,\exp\left(-n\left[\,\sup_{\rho\geq 1}\Big{\{}{{E}_{\mathrm{x}}}\big{(}Q,\rho\big{)}-t{E}_{\mathrm{s}}\big{(}\rho,P_{V}\big{)}\Big{\}}\right]\right),$ (25)
where
$\begin{split}&{E}_{\mathrm{x}}(Q,\rho)\\ &\triangleq\sup_{a(.)}-\rho\sum_{x}Q(x)\log\sum_{\bar{x}}Q(\bar{x})\!\left(e^{-d_{B}(x,\bar{x})}\frac{e^{a(\bar{x})}}{e^{{a(x)}}}\right)^{\frac{1}{\rho}}.\end{split}$ (26)
Corollary 4.
If the codeword distribution in Lemma 5 is chosen as
$Q=\underset{\tilde{Q}}{\operatorname{argmax}}\,\bigg{\{}\sup_{\rho\geq 1}\,{E}_{\mathrm{x}}\big{(}\tilde{Q},\rho\big{)}-t{E}_{\mathrm{s}}\big{(}\rho,P_{V}\big{)}\bigg{\}},$ (27)
then there exists a codebook achieving the following exponent
$E_{G,\mathrm{ex}}^{\mathrm{dual}}(t)$ $\triangleq\sup_{\rho\geq 1}\Big{\{}{{E}_{\mathrm{x}}}(\rho)-t{E}_{\mathrm{s}}\big{(}\rho,P_{V}\big{)}\Big{\}},$ (28)
where ${E}_{\mathrm{x}}(\rho)\triangleq\max_{Q}{E}_{\mathrm{x}}(Q,\rho)$.
The expression in (28) is the dual counterpart of $E_{J,1}^{\mathrm{ex}}(t,P_{V})$ given in (9). This result was previously derived by Gallager in [9, Problem 5.16], where the i.i.d. ensemble was used instead of constant-composition coding.
The result of Corollary 4 cannot be directly compared with Corollary 2, since the concave hull of $E_{\mathrm{x}}^{\prime}(\lambda)$ can, in general, be larger than $E_{\mathrm{x}}(\lambda)$, even though $E_{\mathrm{x}}(\lambda)\geq E_{\mathrm{x}}^{\prime}(\lambda)$ always holds. Thus, whether partitioning can generally improve the achievable expurgated exponent remains an open question, as pointed out in [4, 5]. This contrasts with the case of the random-coding exponent, where it is known that partitioning can strictly improve the exponent [2, 3].